OpenStudy (anonymous):

integrate x^4+2x^2+3/ x^2 dx

OpenStudy (mimi_x3):

\[\int\limits\frac{x^{4}+2x^{2}+3}{x^{2}} dx =>\int\limits\frac{x^{4}}{x^{2}} +\frac{2x^{2}}{x^{2}} +\frac{3}{x^{2}}dx =>\int\limits x^{2} +2+\frac{3}{x^{2}} dx\] Then integrate it..